A man deposits Rs 50000 in the fixed account of a bank which pays interest at the rate of 8% p.a. compounded annually.At the end of each year 5% of the interest earned is deducted as government tax. What sum of money is in his account after 2 years?

Grade:10

1 Answers

Vikas TU
4 years ago
Dear student 
Principle = 50000 
amount after 1 year = (50000* 8 *1/100 ) + 50000 = 54000
deducted amount = 4000*5 /100 = 200 
54000 -200 = 53800
Amount after 2 year = 53800* 8 /100 = 4304 
Amount deducted in 2nd year = 4304 * 5/100 = 215.2 
4304 -215.2 = 4088.8
S total amount after 2 year = 53800 + 4088.8 = 57888.8